# QKFormer: Hierarchical Spiking Transformer using Q-K Attention

### v1

@@ 4.3.1 Visualization
In this part, we visualize the Q-K token attention (Stage 1 and Stage 2 of the QKFormer model) on ImageNet. As shown in Figure 3, $A_{t}$ is the $N*1$ token attention vector, and $X^{\prime}$ is the output of the attention process, which is obtained by the mask operation between matrix $K$ and attention vector $A_{t}$. Specifically, the longitudinal axis denotes the channel index of one head, while the horizontal axis denotes the token index. The #tokens in stage 1 and stage 2 are $56^{2}$ and $28^{2}$, respectively. To facilitate visualization, we choose a continuous segment with a length of 100 extracted from the whole token vector. The visualization shows Q-K attention can lead to high sparsity of spikes.
Figure 3: The visualization of Q-K token attention. The white dot means value 1, while the black one means value 0.
Figure 4: Comparison of memory costs between QKTA and SSA under different token numbers. $N$ is the token number.
4.3.2 Memory Consumption
@@
In this experiment, we compare the memory consumption between QKTA (Formula.10) and SSA (Formula.7) under different token numbers. We calculate the memory consumption of a QKTA and an SSA on a GPU by forwarding the input tensor $(T,B,C,N)$. To facilitate the statistics of the impact of #tokens $N$ on memory consumption, the #channels $C$ is set to 256, and the time step $T$ and batch size $B$ are set to 1. The experiment result is shown in Figure 4. With the increment of #Tokens, SSA consumes much more GPU memory than QKTA, of which the complexity is linear to #Tokens. For example, SSA consumes about 10×10$\times$10 × GPU memory than QKTA when $\sqrt{N}=50$.
4.4 Verification on PEDS Module
In this experiment, we conduct the ablation study for our Patch Embedding module with Deformed Shortcut (PEDS) on CIFAR100 and CIFAR10-DVS datasets.
Table 4: Ablation study results of PEDS on CIFAR100 and CIFAR10-DVS dataset.
Dataset
CIFAR100
CIFAR10-DVS
Spikformer
78.21%
80.9%
Spikformer + PEDS
80.26%
82.2%
We replaced the Spiking Patch Splitting (SPS) module in Spikformer with our PEDS module, while other conditions remain unchanged. The experimental results, shown in Table 4, indicate that our PEDS module brings a great gain in CIFAR100 (+2.05%) and CIFAR10-DVS (+1.3%) compared to the baseline.
4.5 Effects of Time Step
The performance of QKFormer with different simulation time steps of spiking neurons is shown in Table 5. Actually, within a certain range, the larger the time step (T), the better the performance. In addition, our model with T=1 is only 2.64% lower than the model with T=4 on CIFAR100, demonstrating that QKFormer is robust under low latency conditions.
Table 5: The results of QKFormer with different time steps on CIFAR100 classification.
Time step
T=1
T=2
T=4
T=6
QKFormer
78.51%
80.08%
81.15%
81.30%
@@ 5 Conclusion
In this work, we design a novel spike-form Q-K attention considering the properties of SNNs, which can easily model the importance of token or channel dimensions through binary vectors. Q-K attention has linear complexity to #tokens (or #channels) and only adopts two spike-form components: Query ($Q$) and Key ($K$). We design a versatile and powerful Patch Embedding module with a Deformed Shortcut (PEDS), which can improve the performance of spiking transformers greatly. In addition, we develop a hierarchical spiking transformer based on the proposed Q-K attention and PEDS in a direct training way, named QKFormer, which marks the effective exploration of hierarchical spiking representation in Transformer-based SNNs. Extensive experiments show that the proposed model achieves SOTA performance on both static and neuromorphic datasets. Note that QKFormer achieved top-1 accuracy beyond 85% on ImageNet-1k with 4 time steps using the direct training way for the first time. With its powerful performance, we aim for our investigations to instill optimism in the application of SNNs.
Appendix A Appendix
A.1 Experimental Details
A.1.1 Datasets
We evaluate QKFormer on static image classification and neuromorphic classification. The former includes ImageNet-1K Deng et al. [2009], CIFAR10/100 Krizhevsky [2009]. The latter contains CIFAR10-DVS Li et al. [2017] and DVS128 Gesture Amir et al. [2017].
ImageNet-1K is the most typical static image dataset for classification. It contains $1.28$ million images for training and 50k images for validation, with a total of 1,000 categories. CIFAR10/CIFAR100 provides 50, 000 train and 10, 000 test images with 32 × 32 resolution. The difference is that CIFAR10 contains 10 categories for classification. While CIFAR100 contains 100 categories, owning better distinguishing ability for the classification algorithm.
CIFAR10-DVS is an event-based neuromorphic dataset converted from the static image dataset by capturing shifting image samples through the Dynamic Vision Sensor (DVS) camera, which provides 9,000 training samples and 1,000 test samples. DVS128 Gesture is an event-based gesture recognition dataset that contains 1342 samples of 11 hand gesture categories from 29 individuals under 3 illumination conditions, each gesture has an average duration of 6 seconds.
A.1.2 Training Details.
In our experiments, we use 8 NVIDIA Tesla V100 SXM2 32GB GPUs when training models on ImageNet, while 1 GPU is used to train other datasets (CIFAR10, CIFAR100, DVS128 Gesture, CIFAR10-DVS). In direct training SNN models with surrogate function,
$\sigma(x)=\frac{1}{1+\exp(-\alpha x)},$ (18)
we select the Sigmoid function $\sigma(x)$ as the surrogate function with $\alpha=4$ during the backpropagation of direct training in all experiments.
A.2 Q-K Channel Attention
In this experiment, we show three combination strategies of QKCA, QKTA, and SSA. The baseline is our QKFormer, which has 4 blocks, distributed in three stages {1, 1, 2}. The former two blocks use QKTA, and the latter two blocks use SSA. Similarly, we conducted the combination of ”QKCA + SSA” and ”QKTA + QKCA” on CIFAR 100, the experimental results are shown in Table 6.
Table 6: Ablation study results of Q-K attention on CIFAR 100.
Model
QKFormer
QKCA+SSA
QKTA+QKCA
Parameter
6.70 M
6.70 M
6.44 M
CIFAR100
81.15%
81.07%
81.04%
The experimental results show that the three combination strategies have comparable performance on CIFAR 100. Compared with the baseline QKFormer (QKTA + SSA, 6.70 M, 81.15%), both the number of parameters and the classification accuracy of ”QKTA + QKCA” (6.44 M, 81.04%) are reduced. Actually, from our experimental practices, we could choose the one with a larger length of the attention vector for specified network layers when applying QKTA or QKCA to build networks.
A.3 Spike Firing Rates in QKFormer Blocks
In this experiment, we calculate the spike firing rates of QKFormer blocks of the trained QKFormer (64.9M) on the ImageNet-1K test set with the 224 $\times$ 224 input resolution. The average spike firing rates of the QKFormer blocks in Stage1 and Stage2 are shown in Table 7.
Table 7: Spike firing rates in QKFormer blocks. Note that the spike-form $X^{\prime}$ is obtained by column mask operation (token mask) between ${A}_{t}$ and $K$.
QKFormer Block
Stage1
Stage2
QKTA
$Q$
0.0432
0.0231
$K$
0.1784
0.0847
${A}_{t}$
0.3477
0.2655
$X^{\prime}$
0.0832
0.0350
$X^{\prime\prime}$
0.1478
0.0577
SMLP
Layer1
0.0518
0.0246
Layer2
0.2733
0.1869
Table 8: Detailed values of memory consumption of Figure 4 in the main body of this paper.
$\sqrt{N}$
QKTA (M)
SSA (M)
SSA / QKTA
10
0.10
0.14
1.37
20
0.40
1.00
2.53
30
0.89
3.97
4.46
40
1.58
12.19
7.71
50
2.47
26.44
10.70
60
3.56
53.52
15.04
70
4.84
97.64
20.17
80
6.32
162.50
25.70
90
8.18
258.19
31.55
100
10.35
391.77
37.85
110
12.14
570.69
47.01
120
14.23
806.06
56.66
130
16.70
1107.50
66.33
140
20.22
1485.14
73.43
150
22.26
1954.03
87.79
160
26.29
2525.00
96.03
170
28.55
3214.30
112.57
180
32.37
4036.88
124.71
190
36.41
5007.25
137.51
200
40.46
6143.06
151.84
A.4 Supplementary for Memory Consumption in Experiment 4.3
Table 8 shows the detailed values of Figure 4 in the main body of this paper (Experiment 4.3).
A.5 QKFormer with Other Spiking Neurons
Table 9 shows the results of QKFormer with other spiking neuron models: IF Burkitt [2006] and Parametric LIF (PLIF) Fang et al. [2021b]. The time steps of models on CIFAR100 and DVS128 are 4 and 16, respectively.
Table 9: The results of QKFormer with other spiking neuron models on CIFAR100 and DVS128.
Dataset
CIFAR100
DVS128
QKFormer (IF)
80.95%
96.9%
QKFormer (PLIF)
81.12%
98.3%
A.6 Performance Discussion on ImageNet
We show the training loss, test loss, top-1, and top-5 test accuracy of QKFormer (64.96M, 29.08M, 16.47M) on ImageNet-1K in Figure 5. Our QKFormer is a directly trained event-driven SNN on the ImageNet-1K. Under the same experiment conditions without pre-training or extra training data, our QKFormer has surpassed most well-known Transformer-based ANNs with high-power MAC operation: QKFormer (64.96M, 85.65%, SNN) $>$ Swin Transformer(88M, 84.5%, ANN) Liu et al. [2021] $>$ DeiT-B (86M, 83.1%, ANN) Touvron et al. [2021] $>$ ViT (85.9M, 77.9%, ANN) Dosovitskiy et al. [2020].
(a) Training loss
(b) Test loss
(c) Top-1 test accuracy
(d) Top-5 test accuracy
Figure 5: Training loss, test loss, top-1 and top-5 test accuracy of QKFormer on ImageNet-1K. The input resolution of training and testing are 224 $\times$ 224.

### v2

@@ 4.3 Analyses on Q-K Attention
Attention Visualization. In this part, we visualize the Q-K token attention (Stage 1 and Stage 2 of the QKFormer model) on ImageNet. As shown in Figure 3, ${\mathbf{A}}_{t}$ is the $N*1$ token attention vector, and $\mathbf{X}^{\prime}$ is the output of the attention process, which is obtained by the mask operation between matrix $\mathbf{K}$ and attention vector ${\mathbf{A}}_{t}$. Specifically, the longitudinal axis denotes the channel index of one head, while the horizontal axis denotes the token index. The #tokens in stage 1 and stage 2 are $56^{2}$ and $28^{2}$, respectively. To facilitate visualization, we choose a continuous segment with a length of 100 extracted from the whole token vector. The visualization shows Q-K attention can lead to high sparsity of spikes.
@@
Memory Consumption. In this experiment, we compare the memory consumption between QKTA (Formula.6) and SSA (Formula.4) under different token numbers. We calculate the memory consumption of a QKTA and an SSA on a GPU by forwarding the input tensor $(T,B,C,N)$. To facilitate the statistics of the impact of #tokens $N$ on memory consumption, the #channels $C$ is set to 256, and the time step $T$ and batch size $B$ are set to 1. The experiment result is shown in Figure 3. With the increment of #Tokens, SSA consumes much more GPU memory than QKTA, of which the complexity is linear to #Tokens. For example, SSA consumes about 10×10$\times$10 × GPU memory than QKTA when $\sqrt{N}=50$.
Table 4: Spike firing rates in QKFormer blocks.
QKFormer Block Stage1 (fr) Stage2 (fr)
QKTA $\mathbf{Q}$ 0.0432 0.0231
$\mathbf{K}$ 0.1784 0.0847
${\mathbf{A}}_{t}$ 0.3477 0.2655
$\mathbf{X}^{\prime}$ 0.0832 0.0350
$\mathbf{X}^{\prime\prime}$ 0.1478 0.0577
SMLP Layer1 0.0518 0.0246
Layer2 0.2733 0.1869
Spike Firing Rates in QKFormer Blocks. In this experiment, we calculate the spike firing rates of QKFormer blocks of the trained QKFormer (64.9M) on the ImageNet-1K test set with the 224 $\times$ 224 input resolution. The average spike firing rates of the QKFormer blocks in Stage1 and Stage2 are shown in Table 4. Note that the spike-form $\mathbf{X}^{\prime}$ is obtained by column mask operation (token mask) between ${\mathbf{A}}_{t}$ and $\mathbf{K}$. In fact, the summation operation in the Q-K attention causes $\mathbf{Q}$ to become significantly sparser compared to $K$ when the network converges. Specifically, $\mathbf{Q}$ in stage 1 has a fire rate of 0.0432, while $\mathbf{K}$ has 0.1784. After the accumulation operation along $D/h$ of the multi-head QKTA version, the LIF neuron (${\mathbf{A}}_{t}$) exhibits a typical averaged fire rate of 0.3477.
The Variance and Expectation of QKTA. Figure.4 visualize the variance and expectation of QKTA (Formula.15 and 16 in Appendix.6.2) and SSA (Formula.19 and 20 in Appendix.6.2). $N$ is set as 196 and $d$ is set as 64, respectively. We can find SSA has a much larger variance and expectation than QKTA on the whole. For example, the maximum theoretical variance of QKTA is 16, but the maximum theoretical variance of SSA is over 3000. This is the main reason that Q-K attention can discard scaling operations thus reducing power consumption, but SSA can not.
4.4 Ablation Study
SPEDS Module. In this experiment, We replaced the Spiking Patch Splitting (SPS) module in Spikformer with Spiking Patch Embedding with Deformed Shortcut (SPEDS) module, while other conditions remain unchanged. The results (Table 5) show that the SPEDS module is essential to QKFormer on both static and neuromorphic datasets. In addition, the addition of SPEDS to Spikformer leads to great gains in CIFAR100 (+2.05%) and CIFAR10-DVS (+1.30%), which further verified the effectiveness of SPEDS.
Table 5: Ablation studies of SPEDS module.
Model CIFAR100 (Acc) CIFAR10-DVS (Acc)
QKFormer (QKTA + SSA, baseline) 81.15% 84.00%
QKFormer (QKTA + SSA, w/o SPEDS) 80.08% 83.40%
Spikformer (SSA, w/o scaling) 76.95% 79.30%
Spikformer (SSA) 78.21% 80.90%
Spikformer (SSA) + SPEDS 80.26% 82.20%
Table 6: Ablation studies of Q-K Attention.
Model CIFAR100 (Acc, Param) CIFAR10-DVS (Acc, Param)
QKFormer (QKTA + SSA, baseline) 81.15%, 6.74M 84.00%, 1.50M
QKFormer (QKCA + SSA) 81.07%, 6.74M 84.30%, 1.50M
QKFormer (QKTA + QKCA) 81.04%, 6.44M 83.10%, 1.44M
QKFormer (SSA) 81.23%, 6.79M 84.10%, 1.52M
QKFormer (QKCA) 81.00%, 6.44M 80.70%, 1.44M
QKFormer (QKTA) 79.09%, 6.44M 80.70%, 1.44M
Mixed Spiking Attention Integration with Q-K Attention. In this part, we show different integration strategies of QKCA, QKTA, and SSA. The baseline is our QKFormer (QKTA + SSA, 6.70M). The experimental results (Table 6) show that using a single type of Q-K attention (QKTA or QKCA only) in a hierarchical architecture model leads to suboptimal performance. In particular, the performance decline in QKTA is more obvious. While the mixed spiking attention solutions, such as QKFormer(QKTA + QKCA), QKFormer(QKTA + SSA), and QKFormer(QKCA + SSA) can achieve comparable performance to QKFormer(SSA) while requiring fewer parameters and much fewer memory resources (Figure 3). Consequently, the mixed spiking attention solutions are well-suited for larger architectures and more challenging scenarios when considering both computational efficiency and performance.
Table 7: Ablation studies of RC, SN, TS.
Model CIFAR100 (Acc)
QKFormer (baseline) 81.15%
QKFormer (ABA->PA) 81.18%
QKFormer (LIF->IF) 80.95%
QKFormer (LIF->PLIF) 81.12%
QKFormer (T=1) 78.51%
QKFormer (T=2) 80.08%
QKFormer (T=4) 81.15%
QKFormer (T=6) 81.30%
Residual Connection (RC) & Spiking Neuron (SN) & Time Step (TS). The experimental results are shown in Table 7. In this block, we replaced the Activation-Before-Addition (ABA) [28, 11] residual connection of QKFormer with the Pre-activation (PA) [31, 37] way, and the performance slightly improved. In addition, we replaced the LIF spiking neuron with Integrate-and-Fire (IF) and Parametric-Leaky-Integrate-and-Fire (PLIF) [38], which led to slight performance degradation. The accuracy regarding different simulation time steps of QKFormer is shown in the last column. When we increase the simulation time steps, the performance of QKFormer can be further improved. Specifically, QKFormer achieves 81.30 % accuracy on CIFAR 100 when T=6.
@@ 5 Conclusion
In this work, we design a novel spike-form Q-K attention considering the properties of SNNs, which can easily model the importance of token or channel dimensions through binary vectors. Q-K attention has linear complexity to #tokens (or #channels) and only adopts two spike-form components: Query ($\mathbf{Q}$) and Key ($\mathbf{K}$). We design a versatile and powerful Spiking Patch Embedding with Deformed Shortcut (SPEDS), which enhances spiking information transmission and integration, thus improving the performance of spiking transformers. In addition, we develop a hierarchical spiking transformer based on the proposed Q-K attention and SPEDS in a direct training way, named QKFormer, which marks the effective exploration of hierarchical spiking representation in Transformer-based SNNs. Extensive experiments show that the proposed model achieves SOTA performance on both static and neuromorphic datasets. Note that QKFormer achieved top-1 accuracy beyond 85% on ImageNet-1k with 4 time steps using the direct training way for the first time. With its powerful performance, we aim for our investigations to instill optimism in the application of SNNs.
Limitation. Currently, our model is limited to image / DVS classification tasks. We will extend this work to more tasks, such as segmentation, detection, and language tasks, to test the generalizability in the further. In addition, we will explore efficient and high-performance network architectures with fewer time steps based on Q-K attention and other efficient modules, to further reduce the training consumption.